OpenStudy (koury_with_a_k):

I really need help with two physics questions, if anyone's willing to explain them and help. PLEASE! A man uses a lever jack to lift a 2.00 × 10^3 kg car to change a tire. Each stroke of the jack lifts the car a distance of 5.00 × 10^(-3) m. The vertical displacement of the handle is 0.75 m for each stroke. How much total force does he use with each stroke of the jack? A. 3.0 x 10^5 N B. 2.9 x 10^6 N C. 131 N D. 19,600 N

OpenStudy (anonymous):

Let force applied on jack be F, mass of car is m, vertical displacement of car with every stroke is \[\Delta x\], length of handle is l. Then work done on the jack = potential energy gained by car(by its upward progress) so, \[F * l = m*g*\Delta x\] Hence, \[F = m*g*\Delta x /l\] this comes out to be 131N approximately

OpenStudy (koury_with_a_k):

Thank you so much for explaining @zaphodplaysitsafe! I greatly appreciate it. Thanks again!!

OpenStudy (anonymous):

goodluck :)

rebeccaxhawaii (rebeccaxhawaii):

\(\Huge\boxed{\boxed{\boxed{\huge\mathbb{WELCOME~TO~OPENSTUDY} }}}\) Hey i have made this to hopefully answer any questions you may have. But if it doesn't feel free to pm me c: http://openstudy.com/users/rebeccaxhawaii#/updates/56f5205ce4b07a8c82287f75 Good Luck with you studies! \(\Huge\heartsuit\)